OpenStudy (anonymous):

Tasha is planning an expansion of a square flower garden in a city park. If each side of the original garden is increased by 7 m, the new total area of the garden will be 144 m². Find the length of each side of the original garden. * 19 m * 12 m * 5 m * √5 m

OpenStudy (radar):

Let x = original length of side of square garden then, x + 7 will equal the length of the new side for the expanded garden. We know that the Area of a square is equal to the side squared and in this the area is 144 square meters.\[(x+7)^{2}=144 m ^{2}\] Take the square root of both sides getting:\[\sqrt{(x+7)^{2}} =\sqrt{144}\]\[x+7=12m\]x=12-7=5 Original sides = 5 meters Did you follow with understanding?

OpenStudy (anonymous):

yeah i think so

OpenStudy (radar):

If there is a part here you don't understand, I will be more than willing to provide additional help.

OpenStudy (anonymous):

thank you now what if the area was 169m^2

OpenStudy (anonymous):

found the answer nvm

OpenStudy (radar):

Great, and good luck to you in your studies.